OpenStudy (anonymous):

let n be a randomly selected integer from 1 to 20. find the indicated probability (n is prime given that it has 2 digits)

OpenStudy (anonymous):

P(B|A)=P(AandB)/P(A) P(prime|2digits)=P(prime and 2digit)/P(prime) = (4/20) / (11/20) = 4/11